# algorithmic

### v1

@@ 3.1 Utility function design
Figure 3: Utility function contour plots. First, in sub-figure a), we can observe (as done by multiple methods such as Tian et al. [2023a]) that using the average accuracy as a utility does not encourage more claims. In sub-figures b), c) and d), we observe that the utility function encourages the agent to produce as many claims as possible and to maximize the accuracy. It also elegantly ranks samples with different number of claims and accuracy. Furthermore, in subfigure b), we observe that the agent should abstain from a query and obtain a utility of 0 instead of a negative utility if it believes that less than $\rho^{*}=20\%$ of the claims in its best sample are true. Finally, we observe a similar pattern for c) and d), where the target accuracy $\rho^{*}$ is set to 50% and 80% respectively. We also observe the advantages of the utility function ${\mathcal{U}}$ over the average probability empirically in Figure 11.
In order to design a utility function, we assume that 1) a generation $y$ can be broken up into atomic claims $\{c\}_{i=1}^{N}$ as shown in Figure 2, and 2) that each claim can be judged by an oracle ${\mathcal{T}}$ as being factual or not. Previous works, such as [Tian et al., 2023a], have been using the average accuracy of generations to rank them. While this works well to score generations of the same length, it cannot readily be used to score generations of different lengths or abstention, as it does not take into account the number of claims in a generation (see Figure 3 a)). Instead, we design a utility function ${\mathcal{U}}(x,y,\lambda)$ for a query $x$ that gives a utility of 1 for every true claim and a negative utility of $\lambda$ for every false claim. Specifically, we define our utility function as:

### v2

@@ 3.1 Utility function design
Figure 3: Utility function contour plots. First, in sub-figure a), we can observe (as done by multiple methods such as Tian et al. [2023a]) that using the average accuracy as a utility does not encourage more claims. In sub-figures b), c) and d), we observe that the utility function encourages the agent to produce as many claims as possible and to maximize the accuracy. It also elegantly ranks samples with different number of claims and accuracy. Furthermore, in subfigure b), we observe that the agent should abstain from a query and obtain a utility of 0 instead of a negative utility if it believes that less than target accuracy $\rho^{*}=20\%$ of the claims in its best sample are true. Finally, we observe a similar pattern for c) and d), where the target accuracy $\rho^{*}$ is set to 50% and 80% respectively. We also observe the advantages of the utility function ${\mathcal{U}}$ over the average probability empirically in Figure 11.
Limitation of only using accuracy as an utility function. Previous works, such as FActScore [Min et al., 2023] and Finetuning for factuality [Tian et al., 2023a], have been using the average accuracy of generations to rank generations. Where more accurate generation are given a higher score than less accurate ones. While this works well to score generations of the same length, it cannot readily be used to score generations of different lengths or abstention, as it does not take into account the number of claims in a generation (see Figure 3 a)). For example, the model would obtain the same score by providing a single true claim than the score it would obtain by providing 1000 true claims. Therefore, we need to introduce an utility function that take into account the number of claims produced by the model.
Designing an utility function that can handle abstention and generations of different length. Similarly to the FActScore metric (accuracy), we assume that 1) a generation $y$ can be broken up into a set of atomic claims $\{c\}_{i=1}^{N}$ as shown in Figure 2, and 2) that each claim can be judged by an oracle ${\mathcal{T}}$ as being factual or not. Similarly to FActScore the oracle will be a larger LM that has access to Wikpedia while the generating model does not. For query $x$ and answer $y$, our utility function ${\mathcal{U}}(x,y,\lambda)$ gives a utility of 1 for every true claim in $y$ and a negative utility of $\lambda$ for every false claim. Specifically, we define our utility function as:
@@ 4.1 Self-reflection can be used to reduce hallucinations without any external references
FActScore benchmark.
We evaluate the performance of our trained models without any additional finetuning on the FActScore benchmark [Min et al., 2023]. We only benchmark the Llama2 7b chat models on the 183 labelled entities to compare to the results reported by Kang et al. [2023]. In Table 3, we observe that our models strongly outperform every baseline. This is partly due to the limitation of the FActScore metric which, as noted by the authors, measures only precision and not recall. And our models which learn to abstain when unsure can achieve much higher precision.
Model FActScore
Llama 2 7B Dola [Chuang et al., 2024] 36.8
Llama 2 7B Ever + (NRG +SQ) [Kang et al., 2023] 46.7
Llama 2 7B search DPO 85.8
Llama 2 7B ReSearch DPO (Ours) 87.6
Llama 2 7B ReSearch SFT (Ours) 93.0
Table 3: FActScore results. We observe that our models trained on synthetic data generated by ReSearch strongly outperform the baseline on the FActScore benchmark.